SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (03/11/08)

Attendees: Rahmat Rahmat, Mark Tibbets, Rainer Bartoldus, Su Dong

Meeting Information

 Date      :  March 11th, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

    Trigger Configuration was changed on Friday about 3pm.

    There were many error messages recorded by L3 plots for many days, it showed some errors come from SVT or DCH.

    The first bin of GLT object count is still about 1-3 tick(s) away from reference. In our L1 Monitoring plots, we can see that the shift is still consistent within uncertainties

EMT

    Mark is back

DCT

GLT

DQM

    We had a long discussion about request from DQG to prepare the reprocessing.

    Some runs will be marked only bad or good.

    Some runs are in pending, because need Rainer hard work(expected about 1 day)

Online

L3

    Rainer and Silke are working to calculate the luminosity after changes in boost

Trickle injection monitoring

Trigger Configuration

AOB


This page was created by Rahmat; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: Feb 26th, 2008.